OpenStudy (anonymous):

The quadratic equation 2px^2+qx+7=0,where pand q are constants,has two equal roots.Explain p in terms of q.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

say the two equal roots are \(\large \rm \alpha , ~\alpha \)

ganeshie8 (ganeshie8):

\[\large 2px^2 + qx+7=0\] dividing through out \(2p\) gives you \[\large x^2 + \frac{q}{2p}x+\frac{7}{2p}=0\]

ganeshie8 (ganeshie8):

since \(\large ~\alpha, \alpha \) are roots, \(\large (x-\alpha)(x-\alpha)\) is a factor, agree ?

OpenStudy (anonymous):

agree

ganeshie8 (ganeshie8):

you can set them equal : \[\large x^2 + \frac{q}{2p}x+\frac{7}{2p} = (x-\alpha)(x-\alpha)\]

ganeshie8 (ganeshie8):

expand right hand side and compare coefficients both sides

OpenStudy (anonymous):